Given the following data, use exponential smoothing (α=0.2) to develop a demand forecast. Assume the forecast for the initial period is 5. Period Demand 4 5 7 913 8 The exponential smoothing forecast is (round your responses to two decimal places): Period 3 刄 7 Forecast 5.00
Answer #1

Exponential Smoothening:-

Ft = Ft-1 +α(At-1-Ft-1)

In the question, α=0.2, Ft-1 =5

For period 2:-Ft = Ft-1 +α(At-1-Ft-1) =5+0.2(7-5) = 5.4

For period 3:-Ft = Ft-1 +α(At-1-Ft-1) = 5.4 + 0.2(9-5.4) = 6.12

For period 4= Ft = Ft-1 +α(At-1-Ft-1)=6.12+0.2(5-6.12) = 5.9

For period 5 =Ft = Ft-1 +α(At-1-Ft-1) =5.9 + 0.2(9-5.9) =6.52

For period 6 =Ft = Ft-1 +α(At-1-Ft-1) =6.52 + 0.2(13-6.52) =7.82

For period 7 =Ft = Ft-1 +α(At-1-Ft-1) =7.82 + 0.2(8-7.82) =7.86